OpenStudy (anonymous):

I have a problem here..... a=b a^2=ab a^2 - b^2=ab-b^2 (a+b)(a-b)=a(a-b) a+b=a 2a=a 2=1 Which step is wrong?

OpenStudy (chaise):

third line, i think.

OpenStudy (anonymous):

definitely third

OpenStudy (anonymous):

you can subtract a common term from both sides of the equation

OpenStudy (anonymous):

no second is fine mutiplying both sides with a is possible

OpenStudy (anonymous):

no my bad,it's the step 4 to five,you cant eliminate a-b since a=b

OpenStudy (anonymous):

it's the same thing to divide it by zero

OpenStudy (lgbasallote):

^exactly

OpenStudy (shubhamsrg):

yes..what about 4 * 0 = 9* 0 and we have 4=9 !! voila!

OpenStudy (anonymous):

Thanks.....GUYS

OpenStudy (anonymous):

I have got it.....

OpenStudy (maheshmeghwal9):

note carefully:- A CRAZY QUESTION.

OpenStudy (maheshmeghwal9):

but a good one:)

OpenStudy (anonymous):

Why do u think its crazy.......?@maheshmeghwal9

OpenStudy (maheshmeghwal9):

since 2=1. it makes whole math wrong:/

OpenStudy (maheshmeghwal9):

@TareKKowshiK one more thing

OpenStudy (maheshmeghwal9):

(a+b)(a-b)=a(a-b) a+b=a it shouldn't be since (a-b)=0.